Given the b=mpc=.8, and Yf-Ye=400, how much government spending is necessary to close the recessionary gap, and how much tax cut is necessary to get to full employment?

In a depressed economy with a lot of excess capacity and a MPC of 0.9, what effect will a $400 increase in government spending have on equilibrium GDP?

Given, C = 200 + .8Yd , I =300, t =0.25, G = 400, solve the following questions.

What is the multiplier with tax?

what is the equilibrium level of income and employment (Ye)?

What is the amount of disposal income?

What is the consumption level with tax?
Answer #1

Question 1

AD shortfall = Yf - Ye = 400

MPC = 0.8

Calculate Multiplier -

Multiplier = 1/1-MPC = 1/1-0.8 = 1/0.2 = 5

Calculate desired fiscal stimulus -

Desired fiscal stimulus = AD shortfall/Multiplier = 400/5 = 80

Calculate desired tax cut -

Desired tax cut = Desired fiscal stimulus/MPC = 80/0.80 = 100

Thus,

Tax cut of 100 is necessary to get to full employment.